OpenStudy (mellamokatie):

Can someone help me with using Distributive Property to simplify expressions? I know a little on how to do it I just need some help. Thanks.

OpenStudy (kittiwitti1):

\[a(b+c+d...)=ab+ac+ad...\]

OpenStudy (mellamokatie):

Wait I have a problem can I tell you it? Its 5 + 6 (m+1) I think I know how to do it.

OpenStudy (mellamokatie):

Do I first make it look like this: 5+6m(1)?

OpenStudy (twistrose):

no its 9m+9*1

OpenStudy (mellamokatie):

Then do I multiply 6 and 1? And then have 5+7m and then add them and get 12m?

OpenStudy (mellamokatie):

Ohh okay thank you

OpenStudy (mellamokatie):

I still dont understand how you go t 9 though

OpenStudy (kittiwitti1):

\[5+6(m+1)\]Distributive property\[5+(6\times m+6\times1)\]

OpenStudy (mellamokatie):

Okay so I am a little confused the equation switches too 5+(6xm+6+1)?

OpenStudy (kittiwitti1):

If you have variables A, B, and C in equation \(A(B+C)\) you would do this, basically\[A\times B+A\times C\]

OpenStudy (kittiwitti1):

Or A(B) + A(C), remember to do multiplication first :)

OpenStudy (mellamokatie):

Wow that's confusing sorry xD So after it will look like 5(6m+6)?

OpenStudy (mellamokatie):

And would I multiply 5 by 6 and get 30m. After would I add 30m by 6 and get 36m?

OpenStudy (mellamokatie):

@kittiwitti1

OpenStudy (mellamokatie):

@Twistrose can you help me out?

OpenStudy (mellamokatie):

Hey, can you pick up where we left off?

OpenStudy (twistrose):

ok so i go t9 by adding like terms at the beginnning, then distibuting. that

OpenStudy (mellamokatie):

Adding the like terms means 5 and 6 right?

OpenStudy (afro):

you're suppose to do parenthesis first, then like terms 5+6m+6 (Pemdas)

OpenStudy (mellamokatie):

But its distributive property you carry the like terms to to perenteses

OpenStudy (kittiwitti1):

Yeah, you have to do the 6(m+1) part first before anything. Parentheses come before addition.q

OpenStudy (mellamokatie):

Hi kittiwitt1 sorry I just needed some help haha

OpenStudy (afro):

the distributive property counts on what it's near to if the problem was something like this, it would of made more sense to do it like you say (m+1)(6+5)

OpenStudy (kittiwitti1):

5+6(m+1) = 5 + 6m + 6(1) = 5 + 6m + 6 = ?

OpenStudy (afro):

but it's very basic distributive property problem so you're fine 5+6(m+1)= 5+6m+6

OpenStudy (afro):

yea you don't touch the 6m so you just add like terms

OpenStudy (afro):

11+6m last answer

OpenStudy (brainzonly):

Try: 2(4+7) Distribute: 

OpenStudy (kittiwitti1):

@brainzonly is very smart, ask him for help xD I have to finish my Chem homework x_x

OpenStudy (brainzonly):

too many people on this question LOL

OpenStudy (afro):

^

OpenStudy (mellamokatie):

Ugh Chemistry that's hard. Omg this is alot @-@ okay so 5 + (m+6) turns into 5+6m+6 and then I solve that? Ikr but I feel special :DD

OpenStudy (brainzonly):

@Kevin @kittiwitti1 There are certain people that like to exaggerate my talents. :|

OpenStudy (mellamokatie):

LOOL

OpenStudy (mellamokatie):

Thanks for the medal :D

OpenStudy (afro):

lol it's simple algebra you should understand the concept

OpenStudy (brainzonly):

5 + (m+6) could actually be solved this way: 5 + 1(m+6) invisible number :D 5+ 1m +6 1m + 11

OpenStudy (afro):

woah stop brain stop stop

OpenStudy (brainzonly):

? This is easy!!!!!

OpenStudy (afro):

brah's 5+6(m+1) where does that come from

OpenStudy (kittiwitti1):

Brainzzzz slow down,,, lol

OpenStudy (afro):

where does that unknown equation come from chill my mans

OpenStudy (mellamokatie):

I dont understand because first I got 5+6m+6 its really 5 + 1 (m+6)?

OpenStudy (afro):

that's the way mellamokatie you're on the right path now add the like terms (the ones with no letters)

OpenStudy (brainzonly):

WHAT!!! NO!! There is an INViSIBLE NUMBER 1 in front of the (m + 6) 5 + 1(m+6) because 1 times m is just m and 1 times 6 is just 6 so it doesn't distub anything and then 5 + 1m + 6 is 1m + 11 and then that is m + 11

OpenStudy (mellamokatie):

LOOL Afro quit telling me the wrong stuff!!

OpenStudy (kittiwitti1):

He's saying 5 + (m+6) has an "invisible number" because any number/value is also a product of itself times 1\(a=a\times1\)

OpenStudy (brainzonly):

LMAO this is too funny

OpenStudy (afro):

omg dude 5+6(m+1) bro COME ONN, it's not wrong you are correct bro @ katie

OpenStudy (kittiwitti1):

I mean\[a=a\times1\] and Afro is correct

OpenStudy (afro):

 COME ONNNNNNNNN MAN this IS 9TH GRADE STUFF COME ONNNNNN come onnnnnnn